# Moment Tensor Potential and Equivariant Tensor Network Potential with explicit dispersion interactions

### v1

@@ II.9 Assessing the quality of MLIPs
To compare the quality of the fitted MLIPs we evaluated training and validation errors of predicting energies and forces. In addition, we designed two series of tests. In the first series, we evaluated the accuracy of potentials in predicting intermolecular interactions. To that end, we predicted binding curves for a diverse set of dimers and compared them with the ones obtained with ab initio calculations. We calculated the deviation between the position and depth of the minima, as well as the full width of the well at half its depth (referred to as full-width at half-minimum or FWHM) obtained with potentials and ab initio calculations (see Figure 2). In this test, the binding curves were represented as dependencies of energy on the distance between the centers of mass of individual molecules. The second series of tests aimed to assess the accuracy of density and RDF calculated during MD, performed with LAMMPS softwareThompson et al. (2022). We used experimental data as reference.
Figure 2: Error metrics proposed for assessing accuracy of MLIPs in predicting dimer binding curves.
To ensure that the error estimations are statistically meaningful, we conducted each test for an ensemble of five potentials for each type of MLIP model considered, except for calculating density and RDF with ETN+D3, as this type of MLIP is more computationally demanding than others.
@@ III Results and discussion
In this section, we present results of application of the described MLIPs to the modeling of liquid carbon tetrachloride, liquid methane, and liquid toluene. For all systems, we focused on evaluating the accuracy differences between MLIP(l)’s and MLIP+D’s.
@@
We note that for carbon tetrachloride, we performed two test cases. In the first case, we compared the accuracy of MLIP(l)’s and MLIP+D’s when the $R_{\rm cut}$ for all potentials is set to 5.5 Å, which limits the ability of MLIP(l)’s to fully capture the range of dispersion interactions. In the second case, we explore to what extent the accuracy of MLIP(l)’s can be improved by increasing the $R_{\rm cut}$ to 7.5 Å, allowing them to fully capture the range of dispersion interactions (while maintaining a cutoff radius of 5—6 Å for MLIP+D’s). For methane and toluene, we took only the $R_{\rm cut}$ of 7.5 Å for the MLIP(l)’s and of 6 Å for MLIP+D’s.
III.1 CCl4 modeling with $R_{\rm cut}=5.5\text{\AA}$
III.1.1 Datasets generation
To create the pretraining set, we performed AIMD at 298 K for 1 ps in the NVT ensemble. In the subsequent AL process, new configurations were sampled during MD conducted in the NpT ensemble at 293 K and 1 bar. The accumulated training set consisted of 478 configurations, each consisting of 6 CCl4 molecules with a simulation cell size ranging from 10 to 12 Å.
Ab initio simulations were performed using the CP2K package Kühne et al. (2020). We used the Gaussian Plane Wave (GPW) approach Lippert, Hutter, and Parrinello (1997) with QZV2P-GTH-q4 and QZV2P-GTH-q7 basis sets, along with GTH-PBE-q4 and GTH-PBE-q7 pseudopotentials for carbon and chlorine atoms, respectively. The GPW calculation also requires the convergence of the real-space integration grid by adjusting the energy cutoff VandeVondele et al. (2005), which was set to 400 Ry. All calculations were performed at the PBE-D3 level of theory.
III.1.2 Training and validation
For each of the MLIP models, an ensemble of five potentials was trained. The training and validation root mean square errors (RMSEs), averaged over the ensembles, are presented in Figure 3. They indicate an improvement in energy accuracy when moving from MTP to MTP+D2, while the difference between ETN and ETN+D2, ETN+D3 is not clearly distinguishable.
Figure 3: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CCl4 configurations obtained with MLIPs with $R_{\rm cut}=5.5\text{\AA}$.
III.1.3 Binding curves
In this subsection, we predict dimer binding curves for investigating possible sources of training and validation errors, originating from intermolecular interactions. For this test, we considered a diverse set of dimers (Figure 4) to evaluate the precision across a range of interaction strengths and distances (Table 1).
@@ III.1.3 Binding curves
Figure 4: Dimers of CCl4 used for generation of the binding curves. The labels correspond to description of mutual orientation, based on treatment of CCl4 molecule as tetrahedron. F corresponds to “face”, V – “vertex”, E – “edge”, and, finally, s and e correspond to staggered and eclipsed conformations.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -95 -74 -66 -66 -44 -72 -73 -62 -21 -18
rmin, Å 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.3 7.3
Table 1: Depths and positions of binding curves minima for carbon tetrachloride dimers, calculated with PBE-D3 using CP2K.
The mean absolute percentage error (MAPE) and the RMSE averaged over the dimers and ensembles of MLIPs are illustrated in Figure 5(a). The accuracy in the energy at the minimum improves from ETN to ETN+D2 and further to ETN+D3, while for MTP and MTP+D2 the difference is minor. In addition, MLIP+D’s appear to have a slightly higher accuracy in predicting FWHM than MLIP(l)’s.
Figure 5: Errors of predicting the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
MAPE also indicates that predicting the depth and FWHM is more challenging than predicting the position of the minimum. Thus, MLIP+D’s predict FWHM and minima depth with approximately 20-40% error. Furthermore, the precision is not uniform across different dimers (Figure 5(b)) with the highest error corresponding to the VVs and VVe dimers having the weakest interaction and the farthest minimum position (Table 1). According to these tests, MTP shows slightly better accuracy than ETN, while ETN+D2 and ETN+D3 are more precise than MTP+D2. Additionally, ETN+D3 slightly outperforms other MLIP+D’s.
The mean bias error (MBE) (Figure 6(a)) averaged over the ensembles of MLIPs and dimers demonstrates that all the potentials excluding ETN+D3 tend to underestimate the FWHM, especially the MTP and ETN potentials. However, in contrast to MTP, the ETN potential additionally tends to significantly overbind the dimers. This can be demonstrated on the binding curves, averaged over the ensembles of MLIPs (Figure 6(b)) and presented together with a one-$\sigma$ confidence interval.
Figure 6: a) Mean bias error (MBE) on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
The MLIP(l)’s not only exhibit a larger uncertainty but also do not have the correct $1/r^{6}$ asymptotic behavior. The asymptotic behavior slightly improves when dispersion interactions are included via D2, however there is an observable difference compared to the ab initio binding curve, likely due to the difference between D2 $C_{6}$ coefficients used in the dispersion correction of MLIPs and D3 $C_{6}$ coefficients used in ab initio calculations. The best asymptotic behavior is achieved by ETN+D3.
III.1.4 Molecular Dynamics
We proceeded with MD simulations to quantify the impact of differences between MLIPs observed in the binding curves test on density and RDF. For density estimation, we performed 500 ps MD in the NpT ensemble at 293 K and 1 atm with a timestep of 1 fs. The initial simulation cell had a size of 30 Å and consisted of 162 molecules. The resulting density value was averaged over the last 100 ps of the simulation. The dependency of density on time is shown in Figure 7(a), and the MAPE compared to the experimentHien, Ponter, and Peier (1978) is presented in Table 2.
Figure 7: Predicted with the ensembles of MLIPs: a) dependencies of density on time together with one-$\sigma$ confidence intervals; b) average RDFs together with two-$\sigma$ confidence intervals (the larger confidence interval was chosen for illustrative purposes).
In Figure 7(a) we see that the MTP potential does not maintain stable molecular dynamics in the liquid phase, while this issue is resolved by explicit incorporation of dispersion interactions in MTP+D2, resulting in a 25.6% deviation from the experimentally measured value of 1.594 g/cm3. Unlike MTP, the ETN potential is capable of preserving the liquid phase, likely due to the overbinding of molecules, as demonstrated in the dimer binding curves test. Meanwhile, ETN+D2 and ETN+D3 appear to have higher accuracy compared to ETN, achieving a deviation of approximately 22% from the experimental density.
@@ III.1.4 Molecular Dynamics
MAPE 0.991 0.256 0.256 0.216 0.222
$\pm$ 0.001 $\pm$ 0.035 $\pm$ 0.044 $\pm$ 0.009 $\pm$ 0.015
Table 2: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHien, Ponter, and Peier (1978).
To calculate the RDF, we performed MD starting from the simulation cell with the previously converged density, following a 10 ps temperature equilibration in the NVT ensemble at 293 K. The RDF was averaged over 100 ps in the NVE ensemble. The intensities and positions of the maxima of the calculated RDFs for carbon atoms were compared with those obtained experimentallyLowden and Chandler (1974).
In Figure 7(b) the predicted RDF, averaged over the ensemble of MLIPs, is illustrated together with a two-$\sigma$ confidence interval. All MLIPs produce nearly equivalent RDFs (Table 3), at the same time exhibiting a shift compared to the experimental data, likely due to an underestimation of the density.
MTP+D2 ETN ETN+D2 ETN+D3
MAPE(g) 0.041 0.065 0.056 0.056
$\pm$ 0.034 $\pm$ 0.040 $\pm$ 0.053 $\pm$ 0.053
MAPE(r) 0.036 0.031 0.033 0.033
$\pm$ 0.017 $\pm$ 0.019 $\pm$ 0.020 $\pm$ 0.020
Table 3: MAPE and two-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentLowden and Chandler (1974).
In contrast to the accuracy obtained from the binding curve test, ETN+D3 did not outperform ETN+D2 in predicting density and RDF.
III.2 CCl4 modeling with extended $R_{\rm cut}$
In this subsection, we test MLIP(l)’s with the $R_{\rm cut}$ extended to 7.5 Å, while the $R_{\rm cut}$ of MLIP+D’s is set to 6 Å. In this manner, we aim to investigate the ability of MLIP(l)’s to describe both covalent and dispersion interactions simultaneously, which may be challenging because the dispersion interactions are orders of magnitude weaker than intramolecular interactions and are comparable to the magnitude of MLIP fitting errors.
III.2.1 Datasets generation
For ab initio calculations, we used the VASP softwareKresse and Furthmüller (1996), and chose PAW_PBE pseudopotentials, a 900 eV kinetic energy cutoff, and a k-point grid including only gamma-point. All calculations were performed at the PBE-D3 level of theory.
The procedure for generating the training set with total energies and forces was similar to that described earlier. The AIMD of 500 ps was carried out in the NVE ensemble with positions generated by PackmolMartínez et al. (2009) and velocities initialized at 298 K. The MD conditions in the AL process were the same as those used before. At the end of the AL procedure, the accumulated training set comprised 805 configurations, each containing 11 molecules with a cell size ranging from 15 to 17 Å.
III.2.2 Training and validation
Following the previously described procedure, we trained an ensemble of five potentials for each MLIP model. Training and validation errors are presented in Figure 8 and demonstrate the first impact of increasing $R_{\rm cut}$. From the figure we see that the errors of MLIP(l) decreased (by approximately a factor of three for energy and 20% for forces) and reached the level of MLIP+D errors.
Figure 8: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CCl4 configurations obtained with MLIP(l)’s with $R_{\rm cut}=7.5\text{\AA}$ and MLIP+D’s with $R_{\rm cut}=6\text{\AA}$.
III.2.3 Binding curves
For the present test, we used the same set of dimers as was introduced above (Figure 4). We note that the binding curves were recalculated with VASP. The results are presented in Table 4, demonstrating minor changes in interaction energies and minima positions.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -93 -75 -63 -62 -48 -69 -71 -59 -20 -21
rmin 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.2 7.2
Table 4: Depths and positions of binding curves minima for carbon tetrachloride dimers, calculated with PBE-D3 using VASP.
Similarly to the training and validation errors, energy MAPE and RMSE, obtained for the prediction of the binding curves (Figure 9(a)), decreased by approximately 20 and 15%, respectively, compared to the errors obtained for MLIPs with a smaller $R_{\text{cut}}$ However, no significant differences in accuracy were observed between MLIP(l)’s and MLIP+D’s, except for the reduction of errors and the uncertainty of the FWHM estimation.
Figure 9: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
From Figure 9(b) we see that the errors still depend on the mutual orientation of the molecules. Specifically, the energy MAPE for VVs and VVe dimers remains significantly higher than that for others. In addition, MLIP(l)’s now demonstrate smaller energy errors, while possessing higher errors in predicting FWHM on these dimers compared to MLIP+D’s.
Unlike MLIPs with smaller $R_{\rm cut}$, in the present case we did not observe a systematic hardening of the binding curves, as indicated by the MBE (Figure 10(a)). Specifically, only MTP tends to soften the binding curves, while other MLIPs demonstrate a low MBE together with lower MAPE and RMSE (Figure 9 (a)), indicating an accurate prediction of FWHM. The energy error has also been reduced, and most of the MLIPs tend to slightly underestimate the depth of the minima, while ETN continues to show a preference for overbinding.
Figure 10: a) MBE on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
The large errors in predicting FWHM with the MTP potential might result from less accurate asymptotics compared to other MLIPs. This leads to the existence of unphysical minima on the VVe dimer binding curve (Figure 10(b)).
We note that in the present case, ETN+D3 did not exhibit a higher accuracy than ETN+D2, while being more computationally expensive. Moreover, it did not provide more accurate density and RDF estimations when MLIPs with smaller $R_{\rm cut}$ were tested. For these reasons, we do not perform MD simulations for ETN+D3 in the present case.
III.2.4 Molecular Dynamics
The simulation conditions were the same as for the previous MD tests, while the initial configuration size was 30 Å and it consisted of 168 molecules.
Tables 5 and 6 show the errors in density and RDF calculated by different potentials. Similarly to the binding curves, the errors of all MLIPs are close to each other.
MTP MTP+D2 ETN ETN+D2
MAPE 0.151 0.204 0.170 0.177
$\pm$ 0.011 $\pm$ 0.016 $\pm$0.008 $\pm$ 0.010
Table 5: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHien, Ponter, and Peier (1978).
MTP MTP+D2 ETN ETN+D2
MAPE(g) 0.099 0.062 0.118 0.126
$\pm$ 0.075 $\pm$ 0.051 $\pm$ 0.088 $\pm$ 0.094
MAPE(r) 0.028 0.030 0.027 0.029
$\pm$ 0.013 $\pm$ 0.017 $\pm$ 0.015 $\pm$ 0.013
Table 6: MAPE and one-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentLowden and Chandler (1974).
Thus, we demonstrated that it is sufficient to introduce an $R_{\rm cut}$ capturing the range of dispersion interactions for MLIP(l)’s for accurate modeling of liquid carbon tetrachloride, since the difference in accuracy between the MLIP(l) and MLIP+D potentials is minor for all properties calculated excluding FWHM.
However, we want to emphasize that the MLIP+D models still have advantages for the extended-cutoff MLIP(l)’s. Indeed, employing a smaller $R_{\rm cut}$ while incorporating explicit dispersion corrections enhances the computational efficiency of potential fitting and subsequent simulations. This efficiency stems from the reduced computational cost of DFT calculations for smaller configurations, as well as the faster training and evaluation of energies and forces with potentials that have a smaller cutoff radius.
In view of all the results for CCl4, for the other two systems we decided to limit the tests for MLIP+D’s with $R_{\rm cut}=6\text{\AA}$ and MLIP(l)’s with $R_{\rm cut}=7.5\text{\AA}$.
III.3 CH4 modeling
The molecular structure of methane is essentially the same as that of carbon tetrachloride, suggesting the ability of MLIP(l)’s to accurately learn dispersion interactions. However, in liquid methane the dispersion interactions are weaker than in carbon tetrachloride. This raises the question of whether MLIP(l)’s can accurately capture dispersion interactions together with intramolecular interactions when the difference in their magnitudes increases.
III.3.1 Datasets generation
For liquid methane modeling, all ab initio calculations were performed using VASP, with the same settings as those used for CCl4.
AIMD simulations to collect the pretraining set were performed in the NVE ensemble for 500 fs starting from the configuration generated with PackmolMartínez et al. (2009), with velocities initialized at 298 K. During the AL process, MD simulations were performed at 150 K and 10 bar. At the end of the AL process, the dataset accumulated 306 configurations, each comprising 57 molecules with simulation box sizes ranging from 15 to 18 Å.
III.3.2 Training and validation
As in previous cases, for each MLIP model, an ensemble of five potentials was trained, and the RMSEs averaged over the ensembles are presented in Figure 11. The magnitude of the errors is about four times lower for energies and two times lower for forces than that obtained for CCl4 with MLIPs with extended $R_{\rm cut}$. Similarly to CCl4, no significant differences are observed between MLIP(l)’s and MLIP+D’s, excluding thecg improvement of accuracy from ETN to ETN+D2 and further to ETN+D3.
Figure 11: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CH4 configurations.
III.3.3 Binding curves
We constructed binding curves for the same types of dimers as for CCl4 (Figure 4). The ab initio energies and positions of the minima are presented in Table 7. Compared to carbon tetrachloride, CH4 dimers exhibit closer and shallower minima, furthermore, the interaction energies are within a narrower range compared to CCl4 dimers.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -35 -33 -29 -29 -25 -29 -31 -25 -17 -18
rmin 3.7 3.8 4.1 4.1 4.0 4.0 3.9 4.2 4.6 4.6
Table 7: Depths and positions of binding curves minima for methane dimers.
Figure 12(a) shows the binding curve errors, averaged over the ensembles and dimers. There we observe that the precision in energy of MLIP(l)’s and MLIP+D’s is nearly equal, while the ETN family exhibits higher accuracy than the MTP family. However, similar to the prediction of the CCl4 binding curves (Figure 9), MLIP+D’s still have an advantage having considerably lower RMSE and uncertainty in predicting the FWHM.
Figure 12: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
In contrast to CCl4 binding curves, CH4 errors are not strongly dependent on the mutual orientation of the molecules, showing a uniformly high level of accuracy. This may be attributed to the narrower range of interaction energies in CH4 dimers, which does not result in the discrimination of certain dimers, as seen in the case of CCl4.
According to MBE, all MLIPs tend to underestimate the depth of the minima (Figure 13(a)), while both MTP and ETN also show a significant overestimation of the potential width. However, no unphysical minima are observed for MTP, as in CCl4 binding curves (Figure 13(b)). This suggests that all MLIPs are qualitatively correct and reached a close level of accuracy in this test.
Figure 13: a) MBE on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
Similarly to the CCl4 results with the extended cutoff radius of MLIPs, the ETN+D3 does not outperform ETN+D2 and thus will not be used for further simulations due to being more computationally expensive.
III.3.4 Molecular Dynamics
To predict density, a 30 Å cubic cell, consisting of 456 molecules, was equilibrated for 300 ps in the NpT ensemble, at $T=110$ K and $p=20.75$ bar. The resulting density value was averaged over the last 100 ps and compared to the experimental valueHarris and Trappeniers (1980) of 0.4256 g/cm3. The calculated MAPE is presented in Table 8. The errors obtained for all MLIPs are small and close in magnitude.
MTP MTP+D2 ETN ETN+D2
MAPE 0.033 0.045 0.035 0.043

### v2

@@ II.9 Assessing the quality of MLIPs
To compare the quality of the fitted MLIPs we evaluated training and validation errors of predicting energies and forces. In addition, we designed two series of tests. In the first series, we evaluated the accuracy of potentials in predicting intermolecular interactions. To that end, we predicted binding curves for a diverse set of dimers and compared them with the ones obtained with ab initio calculations. We calculated the deviation between the position and depth of the minima, as well as the full width of the well at half its depth (referred to as full-width at half-minimum or FWHM) obtained with potentials and ab initio calculations (see supplementary material for details). In this test, the binding curves were represented as dependencies of the interaction energy on the distance between the centers of mass of individual molecules. The second series of tests aimed to assess the accuracy of density and RDF calculated during MD, performed with LAMMPS package Thompson et al. (2022). We used experimental data as reference.
For evaluating the errors, we used the root mean square error (RMSE) and mean absolute percentage error (MAPE):
${\rm RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(x_{i}-\hat{x}_{i})^{2}}{N}},$ (22)
${\rm MAPE}=\frac{\sum_{i=1}^{N}|x_{i}-\hat{x}_{i}|}{N},$ (23)
where $N$ is the number of samples, and $x_{i}$ and $\hat{x}_{i}$ are the predicted and true values, respectively.
To ensure that error estimations are statistically meaningful, we conducted each test for an ensemble of five potentials for each type of MLIP model considered, except for calculating density and RDF with ETN+D3, as this type of MLIP is more computationally demanding than others.
@@ III Results and discussion
In this section, we present results of application of the described MLIPs to the modeling of liquid carbon tetrachloride, liquid methane, and liquid toluene. For all systems, we focus on evaluating the accuracy differences between MLIP(l)’s and MLIP+D’s.
@@
We note that for carbon tetrachloride, we performed two test cases. In the first case, we compared the accuracy of MLIP(l)’s and MLIP+D’s when the $R_{\rm cut}$ for all potentials is set to 5.5 Å, which limits the ability of MLIP(l)’s to fully capture the range of dispersion interactions. In the second case, we explored to what extent the accuracy of MLIP(l)’s can be improved by increasing the $R_{\rm cut}$ to 7.5 Å, allowing them to capture the major part of long-range interactions (while maintaining a cutoff radius of 5–6 Å for MLIP+D’s). For methane and toluene, we took only the $R_{\rm cut}$ of 7.5 Å for the MLIP(l)’s and of 6 Å for MLIP+D’s.
III.1 Computational details
For the generation of the training set, we employed the AL scheme described above (Figure 1). The simulation setup for AIMD and AL-MD, details of the ab initio calculations, and sizes of the obtained training sets are presented in Table 1.
ab initio package level of theory basis set pseudo-potential cutoff,
k-points
number of
molecules
in cell
cell size, Å MD training
set size
[…]
QZV2P-GTH-q4,
QZV2P-GTH-q7
GTH-PBE-q4,
GTH-PBE-q7
energy cutoff VandeVondele et al. (2005): 400 Ry; $\Gamma$-point 6 10-12 AIMD: at 298 K for 1 ps in the NVT ensemble;
AL-MD: NpT ensemble at 293 K and 1 bar
478
[…]
$\Gamma$-point
11 15-17 AIMD: 500 ps was carried out in the NVE, velocities initialized at 298 K;
AL-MD: NpT ensemble at 293 K and 1 bar
850
CH4 VASP PBE-D3 PW PAW_PBE 900 eV
$\Gamma$-point
57 15-18 AIMD: NVE ensemble for 500 fs, velocities initialized at 298 K;
AL-MD: NpT ensemble at 150 K and 10 bar.
306
Toluene FHI-aims Blum et al. (2009) PBE-TS Numerically tabulated Atom-centered Orbitals Blum et al. (2009), intermediate basis set all electrons 2x2x2 k-point grid 13 15-18 AIMD: NVT ensemble for 400 fs, at 298 K;
AL-MD: NpT ensemble at 298 K and 1 atm
520
Table 1: Computational details on the generation of the training sets.
In all ab initio calculations, we employed dispersion-corrected DFT at the PBE-D3 or PBE-TS level of theory. The employed D3 version includes the Becke-Johnson damping function and omits the three-body (ATM) term. This means that we do not account for many-body effects, which affect the accuracy of the ab initio method and, consequently, the trained potentials. However, since dispersion in ab initio methods is described only by two-body potentials, this approach is consistent with our use of two-body D2 and D3 terms in the MLIP+D framework. However, if MLIP+Dn models were fitted to reference data that included many-body effects, incorporating the ATM term in the D3 term should be expected to enhance the accuracy.
The chosen ab initio models described above were also employed for the calculation of the dimer binding curves, used to test the resulting MLIPs. In addition to testing MLIPs on the binding curves, we also used them during MD simulations to predict densities and RDFs. To determine the densities, we equilibrated the simulation cells in the NpT ensemble until convergence was achieved, and then averaged the density over 100 ps of the simulation. The simulation details of the density calculations are given in Table 2. For calculating RDFs, the simulations started from the configuration with converged density. Then, the system was equilibrated in the NVT ensemble at the target temperature during 10 ps. Subsequently, the ensemble was changed to NVE, and the RDF was averaged over 100 ps. A timestep of 1 fs was used in all simulations. For CCl4 and toluene, the RDF was obtained under the same conditions as the density, while for CH4 the temperature was set to 92 K and the pressure was set to 0.13 atm to match the conditions of the available experimental data.
number of molecules Density simulation
trajectory length, ps T, K p, atm
[…]
CH4 456 300 110 20.48
Toluene 104 200 298 1
Table 2: Computational details on density calculations.
III.2 Training and validation errors
For all the test cases, ensembles of five potentials were trained for each MLIP type. The training and validation RMSEs, averaged over the ensembles, are presented in Figure 2. For MLIPs with a cutoff of 5.5 Å, we observe a reduction in the energy RMSEs moving from MTP to MTP+D2, while the effect of dispersion correction for ETN-based potentials is less significant and observed only for validation energy RMSEs. When extending the cutoff radius of MLIP(l)’s from 5.5 Å to 7.5 Å, the errors decreased by approximately a factor of three for energy, matching the accuracy of MLIP+D’s. This trend also holds for CH4, where errors are approximately four times smaller for energies and two times lower for forces than those obtained for CCl4 with MLIPs with extended $R_{\rm cut}$. However, the MLIPs fitted on toluene, even when the cutoff radius of MLIP(l)’s is extended, were not able to reach the accuracy level of MLIP+D’s.
[…]
III.3 Binding curves
[…]
@@ III.3 Binding curves
Figure 3: a) Dimers of CCl4 and CH4 used for generation of the binding curves. The labels correspond to description of mutual orientation, based on treatment of CCl4 molecule as tetrahedron. F corresponds to “face”, V – “vertex”, E – “edge”, and, finally, s and e correspond to staggered and eclipsed conformations; b) dimers of toluene used for constructing binding curves and minima depths and positions. The labels denote mutual orientations of molecules: "P" – parallel, "T" – T-shaped, "IP" – in plane.
The minima depth and position on the binding curves for CCl4, CH4 and toluene dimers obtained with ab initio calculations are presented in Table 3. The binding curves were calculated with the same ab initio package that was used to generate the corresponding training set. Compared to carbon tetrachloride, CH4 dimers exhibit closer and shallower minima. In contrast, toluene dimers possess the widest range of minima depths and positions among the studied systems, suggesting that toluene will be the most challenging molecule to model for MLIPs.
$F$s FFe FVs FVe $E$s EEe EF EV $V$s VVe
[…]
r 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.3 7.3
[…]
r 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.2 7.2
CH4 E -35 -33 -29 -29 -25 -29 -31 -25 -17 -18
r 3.7 3.8 4.1 4.1 4.0 4.0 3.9 4.2 4.6 4.6
Toluene P-1 P-2 P-3 P-4 T-1 T-2 T-3 IP-1 IP-2 IP-3
E -173 -172 -134 -115 -123 -139 -104 -26 -46 -57
r 3.8 4.2 5.1 5.8 6.0 5.6 5.9 8.2 8.3 8.4
[…]
The calculated RMSEs for the energy minimum ($E_{\rm min}$), the position of its minimum ($r_{\rm min}$), and the FWHM are shown in Figure 4 (a). We observed that explicit incorporation of the dispersion correction significantly affects the FWHM RMSEs when $R_{\rm cut}$ is set to 5.5 Å for MLIP(l)’s, and it also significantly improves the accuracy of $E_{\rm min}$ for the ETN model (Figure 4 (a)). The effect of incorporating the dispersion correction corresponds to approximately 50% decrease in the FWHM RMSE for MTPs-based potentials and to 50% decrease in the energy RMSE for the ETN-based potentials. However, RMSE changes of the same order of magnitude are observed when extending the MLIP(l)’s cutoff radius to 7.5 Å (Figure 4 (a)). In particular, the ETN energy RMSE and MTP FWHM RMSE decrease by about 50%, reaching the accuracy of MLIP+D’s. Accordingly, for the other systems, since the MLIP(l)’s cutoff radii were extended, the difference between MLIP(l)’s and MLIP+D’s is not significant in any property excluding FWHM in CH4 where MLIP+D’s still show higher accuracy than MLIP(l)’s. At the same time, the RMSEs for CH4 are much smaller than for CCl4, while for toluene they are significantly higher than in other test cases, including CCl4 modeling with MLIPs when $R_{\rm{cut}}$ is set to 5.5 Å. Additionally, we note that predicting the depth and FWHM is more challenging than predicting the position of the minimum, as it is less sensitive to the incorporation of dispersion correction or increasing $R_{\rm cut}$.
[…]
Since we observed the significant impact of dispersion correction in modeling of CCl4 with MLIPs when $R_{\rm cut}$ is set to 5.5 Å, we quantified this impact in absolute values of the energy minima and FWHMs (see supplementary material for absolute values of the minima energy and FWHM for other test cases). As shown in Figure 4 (b), the high energy RMSE of ETN corresponds to a significant overbinding of the majority of dimers, which was mitigated by the explicit incorporation of the dispersion correction in ETN+D2 and ETN+D3. At the same time, the difference between the energies predicted by MTP and MTP+D2 is not significant. We also note that all the potentials, excluding ETN+D3, tend to underestimate FWHM, especially the MTP and ETN potentials.
Figure 5: Toluene dimers binding curves predicted by MLIPs compared to ab initio calculated.
In addition to the findings related to the energy minima and the potential well width, we made a further observation concerning the energy dependence on the intermolecular separation. In Figure 5 we demonstrate the unphysical oscillations in the toluene dimer binding curves predicted by MLIP(l)’s with $R_{\rm cut}=7.5$ Å and MLIP+D’s with $R_{\rm cut}=6$ Å. This behavior is observed for all the MLIPs, but is more severe in the case of the ETN-based ones. These oscillations are weaker for MLIP+D’s compared to MLIP(l)’s. This observation explains the difference in MLIP(l)’s and MLIP+D’s training and validation energy errors, which could not be explained from the energy minima, position, and FHWM RMSEs, as they did not change substantially with the incorporation of dispersion correction. Similarly, such oscillations are observed for CCl4 VVe dimer binding curve predicted by MTP with the extended $R_{\rm cut}$ (see supplementary material for details), while MTP+D2 preserved more correct asymptotics. We note that all binding curves for CCl4 and CH4 are also given in supplementary material.
Finally, the overall improvement in binding curve predictions achieved by incorporation of long-range interactions is both intuitive and well-documented in literature for MLIPs that account for electrostatic interactions Grisafi, Nigam, and Ceriotti (2021); Yue et al. (2021).
III.4 Densities
Next, we calculated the liquid densities for all considered cases. Similar to the binding curve tests, we focus on the impact of adding the dispersion correction and, for CCl4, on extending the cutoff radius for MLIP(l)’s.
All density values were obtained by averaging over 100 ps of the NpT MD trajectories after the density converged. The details of the simulation are given in Table 2. Figure 6 shows the predicted and experimental densities, and Table 4 contains the corresponding MAPEs for all investigated systems. Similarly to the binding curve tests, only CCl4 was modeled with MLIP(l)’s and MLIP+D’s using two sets of cutoff radii.
[…]
@@ III.4 Densities
CCl4, (a) 0.991 0.26 0.26 0.216 0.22
$\pm$ 0.001 $\pm$ 0.04 $\pm$ 0.04 $\pm$ 0.009 $\pm$ 0.02
CCl4, (b) 0.156 0.20 0.170 0.18 -
$\pm$ 0.004 $\pm$ 0.02 $\pm$0.008 $\pm$ 0.01
CH4 0.033 0.045 0.035 0.043 -